If the learning rate is 0.90, and the first 5 units had an average cost of $40, what is the cumulative average cost for the first 100 units?

Enter your answer to two decimal places, with no dollar sign.
Answer #1

Learing equation formula c Daxo ( are of firet cost with b log (LP) = log (0.9); - - - 0.152 log(2) log(2) = a (51-0.152 40 =